OpenStudy (anonymous):

1. Solve for r. A=p + prt 2. Solve for m. Y=mx + b

OpenStudy (anonymous):

For both problems, all you have to do is get r (first problem) or m (second problem) alone on one side of the equation.

OpenStudy (anonymous):

Well I did this and apparently my answers were incorrect. For solving r I came up with a-p/pt=r And for solving m I came up with y-b/x=m

OpenStudy (anonymous):

For the second problem: Y=mx +b ----> Subtract b from both sides you get Y-b=mx ------> divide by x from both sides you get (Y-b)/x=m Just remember that whatever you do to one side of the equation, you have to do to the other side in order to keep both sides equal to each other.

OpenStudy (anonymous):

did you include parentheses?

OpenStudy (anonymous):

No i did not are you suppose to? And for solving r equation do you need parenthesis around a-p

OpenStudy (anonymous):

yes. Parentheses indicate which operation is performed first. For example in the second problem the answer is (Y-b)/x=m. The parentheses indicate that Y-b is done first before dividing by x. Otherwise, your answer: Y-b/x=m. Indicates that you first divide b by x (b/x) and then subtract that result from Y. If you substitute numbers for Y, b, and x, you'll see that the result is different for m.

OpenStudy (anonymous):

Okay thank youu !! & I have another question... lol So i have this word problem and I figured it out but the teacher seems to think this is incorrect. Would you care to see..? The perimeter of a triangle is 28ft. The medium side is 4ft longer than the shortest side,while the longest side is twice as long as the shortest side. What are the lengths of the three sides? A+B+C=P 2c+(4c+c)+c=28 2c+4+c+c=28 4c+4=28 4c=24 divide by 4 c=6 a=12 b=10

OpenStudy (anonymous):

Your answer is correct. However, your second line is not written correctly. You wrote: 2c+(4c+c)+c=28 It should be: 2c+(4+c)+c=28 (this is the same as your third line). Other than that, I don't see anything wrong with it as long as you labeled C as the shortest side, B as the medium side, and A as the longest side.